### v1

@@ 4.2 Medical Visual Mamba
Precise 3D segmentation outcomes can alleviate physicians’ diagnostic workloads in disease management. SegMamba xing2024segmamba, a cutting-edge architecture, is the first method to utilize Mamba specifically for accurate 3D segmentation in medical imaging. It introduced a tri-orientated Mamba (ToM) module for modeling 3D features from three directions and a gated spatial convolution (GSC) module to enhance spatial feature representation before each ToM module. Similarly employing a U-shaped architecture, LightM-UNet liao2024lightm harnesses the Residual Vision Mamba Layer exclusively in a solely Mamba approach for extracting deep semantic features and modeling extensive spatial dependencies within a lightweight framework. Moreover, both LMa-UNet wang2024large and T-Mamba hao2024t built upon the foundation of SegMamba, with improvements made to the Mamba block. A notable aspect of LMa-UNet wang2024large is its use of large windows, which outperforms small kernel-based CNNs and small window-based Transformers in local spatial modeling, while T-Mamba hao2024t develops a gate selection unit to adaptively combine two features in the spatial domain with one feature in the frequency domain, marking the first instance of incorporating frequency-based features into the vision mamba framework. The issue of long-term temporal dependency in video scenarios has also been addressed by developing a generic Video Vision Mamba-based framework named Vivim yang2024vivim. Using the specifically engineered Temporal Mamba Block, this model effectively compresses long-term spatiotemporal data into sequences of different scales.
@@
For image registration task, MambaMorph guo2024mambamorph introduces a groundbreaking multi-modality deformable registration framework that enhances medical image analysis by combining a Mamba-based registration module with an advanced feature extractor for efficient spatial correspondence and feature learning. The VMambaMorph  li20243dmambacomplete has further enhanced its VMamba-based block by incorporating a 2D cross-scan module, redesigned to process 3D volumetric features efficiently.
@@
In other domains, the Contrastive Masked Vim Autoencoder (CMViM) yang2024cmvim tackles Alzheimer’s disease (AD) classification by incorporating Vision Mamba (ViM) into a masked autoencoder for 3D multi-modal data reconstruction. For endoscopy skill evaluation, a low-cost motion-guided dual-camera tracker zhang2024motion provides reliable endoscope tip feedback, and a Mamba-based motion-guided prediction head (MMH) merges visual tracking with historical motion data using a state space model.
@@ 4.2.3 Challenge
Infrared Small Target Detection MIM-ISTD  chen2024mim Mamba-in-Mamba architecture
@@
Classification RSMamba  chen2024rsmamba multi-path activation -
HSIMamba  yang2024hsimamba process data bidirectionally Memory 136.53
Image dense prediction RS-Mamba  zhao2024rs omnidirectional selective scan -
Change detection ChangeMamba  chen2024changemamba cross-scan mechanism -
Semantic segmentation RS3Mamba  ma2024rs3mamba dual-branch network
@@
Samba  zhu2024samba encoder-decoder architecture
@@ 4.3 Remote Sensing Image
Drawing inspiration from TNT, Chen et al. introduced a new Mamba-in-Mamba (MiM-ISTD)  chen2024mim architecture to enhance the efficiency of infrared small target detection. In this approach, local patches are considered "visual sentences," while Outer Mamba is utilized to extract global information. Moreover, for remote sensing image classification, RSMamba  chen2024rsmamba features a dynamic multi-path activation mechanism to improve Mamba’s capability in handling non-causal data. RS-Mamba  zhao2024rs is adept at handling very-high-resolution (VHR) remote sensing images for dense prediction tasks, utilizing an omnidirectional selective scan module to model images from various angles comprehensively. Classification of hyperspectral images is difficult in remote sensing research due to their complex, high-dimensional data.HSIMamba yang2024hsimamba is designed with a module dedicated to spatial analysis, which includes multiple spectral bands and three-dimensional spatial structures to take advantage of the rich multidimensional nature of the hyperspectral data and to enhance the feature representation capability using linear transformations and activation functions. In addition, HSIMamba employs a bi-directional processing mechanism that improves the model’s ability to represent and utilize spectral information through forward and backward spectral dependency capture. Pan-Mamba  he2024pan also provides an innovative network in the pansharpening domain and customizes two crucial components, channel swapping Mamba and cross-modal Mamba, both carefully crafted for efficient cross-modal information exchange and fusion.ChangeMamba chen2024changemamba explores for the first time the potential of the Mamba architecture for remote sensing change detection (CD) tasks. The MambaBCD, MambaSCD, and MambaBDA network frameworks are designed for Binary Change Detection (BCD), Semantic Change Detection (SCD), and Building Damage Assessment (BDA) tasks, and three spatio-temporal relationship modeling mechanisms are proposed to learn spatio-temporal features thoroughly.ChangeMamba utilizes selective state-space modeling to capture long-range dependent features and maintains linear computational complexity while providing Visual Mamba architecture to learn global spatial context information. Semantic segmentation of remotely sensed images is a fundamental task in geoscientific research.RS3Mamba ma2024rs3mamba is a novel two-branch network for semantic segmentation of remotely sensed images. The network incorporates visual state space (VSS) models, especially the Mamba architecture, to improve long-range relational modeling capabilities. In addition, a Co-Completion Module (CCM) is proposed for feature fusion. Experimental results show that RS3Mamba has significant advantages over CNN and transformer-based approaches.Samba zhu2024samba is a novel semantic segmentation framework specialized for high-resolution remote sensing images with encoder-decoder architecture. Samba blocks act as encoders to extract multilevel semantic information, and Mamba blocks utilize state space models to capture global semantic information with linear computational complexity.

### v2

@@ 4.2 Medical Visual Mamba
Precise 3D segmentation outcomes can alleviate physicians’ diagnostic workloads in disease management. SegMamba [98], a cutting-edge architecture, is the first method to utilize Mamba specifically for accurate 3D segmentation in medical imaging. It introduced a tri-orientated Mamba (ToM) module for modeling 3D features from three directions and a gated spatial convolution (GSC) module to enhance spatial feature representation before each ToM module. Similarly employing a U-shaped architecture, LightM-UNet [97] harnesses the Residual Vision Mamba Layer exclusively in a solely Mamba approach for extracting deep semantic features and modeling extensive spatial dependencies within a lightweight framework. Moreover, both LMa-UNet [45] and T-Mamba [99] built upon the foundation of SegMamba, with improvements made to the Mamba block. A notable aspect of LMa-UNet [45] is its use of large windows, which outperforms small kernel-based CNNs and small window-based Transformers in local spatial modeling, while T-Mamba [99] develops a gate selection unit to adaptively combine two features in the spatial domain with one feature in the frequency domain, marking the first instance of incorporating frequency-based features into the vision mamba framework. The issue of long-term temporal dependency in video scenarios has also been addressed by developing a generic Video Vision Mamba-based framework named Vivim [38]. Using the specifically engineered Temporal Mamba Block, this model effectively compresses long-term spatiotemporal data into sequences of different scales.
@@
For image registration task, MambaMorph [102] introduces a groundbreaking multi-modality deformable registration framework that enhances medical image analysis by combining a Mamba-based registration module with an advanced feature extractor for efficient spatial correspondence and feature learning. The VMambaMorph  [79] has further enhanced its VMamba-based block by incorporating a 2D cross-scan module, redesigned to process 3D volumetric features efficiently.
@@
In other domains, the Contrastive Masked Vim Autoencoder (CMViM) [100] tackles Alzheimer’s disease (AD) classification by incorporating Vision Mamba (ViM) into a masked autoencoder for 3D multi-modal data reconstruction. For endoscopy skill evaluation, a low-cost motion-guided dual-camera tracker [37] provides reliable endoscope tip feedback, and a Mamba-based motion-guided prediction head (MMH) merges visual tracking with historical motion data using a state space model.
@@ 4.2.3 Challenge
Infrared Small Target Detection MIM-ISTD  [66] Mamba-in-Mamba architecture
@@
Classification RSMamba  [39] multi-path activation -
HSIMamba  [104] process data bidirectionally Memory 136.53
Image dense prediction RS-Mamba  [69] omnidirectional selective scan -
Change detection ChangeMamba  [54] cross-scan mechanism -
Semantic segmentation RS3Mamba  [67] dual-branch network
@@
Samba  [105] encoder-decoder architecture
@@ 4.3 Remote Sensing Image
Drawing inspiration from TNT, Chen et al. introduced a new Mamba-in-Mamba (MiM-ISTD)  [66] architecture to enhance the efficiency of infrared small target detection. In this approach, local patches are considered "visual sentences," while Outer Mamba is utilized to extract global information. Moreover, for remote sensing image classification, RSMamba  [39] features a dynamic multi-path activation mechanism to improve Mamba’s capability in handling non-causal data. RS-Mamba  [69] is adept at handling very-high-resolution (VHR) remote sensing images for dense prediction tasks, utilizing an omnidirectional selective scan module to model images from various angles comprehensively. Classification of hyperspectral images is difficult in remote sensing research due to their complex, high-dimensional data.HSIMamba [104] is designed with a module dedicated to spatial analysis, which includes multiple spectral bands and three-dimensional spatial structures to take advantage of the rich multidimensional nature of the hyperspectral data and to enhance the feature representation capability using linear transformations and activation functions. In addition, HSIMamba employs a bi-directional processing mechanism that improves the model’s ability to represent and utilize spectral information through forward and backward spectral dependency capture. Pan-Mamba  [103] also provides an innovative network in the pansharpening domain and customizes two crucial components, channel swapping Mamba and cross-modal Mamba, both carefully crafted for efficient cross-modal information exchange and fusion.ChangeMamba [54] explores for the first time the potential of the Mamba architecture for remote sensing change detection (CD) tasks. The MambaBCD, MambaSCD, and MambaBDA network frameworks are designed for Binary Change Detection (BCD), Semantic Change Detection (SCD), and Building Damage Assessment (BDA) tasks, and three spatio-temporal relationship modeling mechanisms are proposed to learn spatio-temporal features thoroughly.ChangeMamba utilizes selective state-space modeling to capture long-range dependent features and maintains linear computational complexity while providing Visual Mamba architecture to learn global spatial context information. Semantic segmentation of remotely sensed images is a fundamental task in geoscientific research.RS3Mamba [67] is a novel two-branch network for semantic segmentation of remotely sensed images. The network incorporates visual state space (VSS) models, especially the Mamba architecture, to improve long-range relational modeling capabilities. In addition, a Co-Completion Module (CCM) is proposed for feature fusion. Experimental results show that RS3Mamba has significant advantages over CNN and transformer-based approaches.Samba [105] is a novel semantic segmentation framework specialized for high-resolution remote sensing images with encoder-decoder architecture. Samba blocks act as encoders to extract multilevel semantic information, and Mamba blocks utilize state space models to capture global semantic information with linear computational complexity.
@@ 5.0.1 Acknowledgements
The authors have no competing interests to declare relevant to this article’s content.